OpenStudy (anonymous):

you have checked out b books from the library,where b is a number greater than 5.you plan to read three of the books and return them at the end of this week. in how many ways can you pick the books you will read,first,second,and third? a.b^3 ways b. (b^3-3b^2+2b)ways c.(b^3-9b^2+27b-27)ways d.(5b-3)

OpenStudy (kropot72):

The number of permutations is bP3: \[bP3=\frac{b!}{(b - 3)!}=b(b - 1)(b - 2)=You\ can\ finish\ now.\]